Marquis Company estimates that annual manufacturing overhead costs will be $801,960. Estimated annual operating activity bases are direct labor cost $489,000, direct labor hours 48,900, and machine hours 97,800.

Compute the predetermined overhead rate for each activity base. (Round answers to 2 decimal places, e.g. 10.50% or 10.50.)

Overhead rate per direct labor cost

enter percentages rounded to 2 decimal places

%
Overhead rate per direct labor hour

$enter a dollar amount rounded to 2 decimal places

Overhead rate per machine hour

$enter a dollar amount rounded to 2 decimal places
Answer #1

Calculation of required overhead rates:-   

a) Overhead rate per direct labor cost = (Annual manufacturing overhead costs/Annual Direct labor cost)*100

= ($801,960/$489,000)*100 = 164.00 %

Therefore overhead rate per direct labor cost is 164%.

b) Overhead rate per direct labor hour = Annual manufacturing overhead costs/Annual Direct labor hours

= $801,960/48,900 = $16.40 per direct labor hour

c) Overhead rate per machine hour = Annual manufacturing overhead costs/Annual machine hours

= $801,960/97,800 = $8.20 per machine hour